40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 548     Entry time: Fri Jun 20 02:20:33 2008
Author: Yoichi 
Type: Update 
Category:  
Subject: PMC transmittance degradation 
The PMC transmitted light power has been degrading constantly for last two weeks (see the attachment 1).
I went down to 2.55V.
The output of the MOPA is constant during this period. More strangely, the reflected power from the PMC is also constant.
One possible explanation is the contamination of the PMC mirrors. But I don't know why it started two weeks ago.

I tweaked the alignment of PMC and was able to recover the transmitted power to above 2.7V (attachment 2).
I will keep eye on this issue.
Attachment 1: pmc_trans_long_trend.pdf  1.226 MB  | Hide | Hide all
pmc_trans_long_trend.pdf
Attachment 2: pmc_trans_improvement.pdf  83 kB  | Hide | Hide all
pmc_trans_improvement.pdf
ELOG V3.1.3-